The elastic limit of steel is 8 × 108 N m-2 and its Young modulus 2 × 1011 N m-2. Find the maximum elongation of a half-meter steel wire that can be given without exceeding the elastic limit.

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. y = F/A L/∆L ⇒ ∆L = FL/AY